# Quantum Quasinormal Mode Theory for Dissipative Nano-Optics and Magnetodielectric Cavity Quantum Electrodynamics

### v2

@@ IV Effective Lindblad master equation
Figure 2: Schematic representation of the quantum description of an electric/magnetic dipolar QE, modeled as a TLS, interacting with an electromagnetic resonator using PMLs. (a) Most general scenario, in which the resonator could have electric and magnetic gain components in distinct spatial or spectral regions (e.g., a pair of coupled, spherical resonators composed of different materials). In this case, every eigenmode of the resonator (both QNMs and PML modes) is described by four bosonic operators $\hat{a}_{n,\lambda,X}$ with $\lambda=\mathrm{e,m}$ and $X=\mathrm{L,G}$, corresponding to differently symmetrized modes according to Eqs. (48) and (49). The presence of gain/loss in the system necessitates a symmetrization and the consequent coupling between the symmetrized modes described by $\chi_{nn^{\prime}}^{\lambda,X,+}$ in the effective Hamiltonian of Eq. (65). The losses of the system are described by Lindblad dissipators proportional to $\tilde{\chi}_{nn^{\prime}}^{\lambda,\mathrm{L},-}$ and the gain components of the PMLs by incoherent pumping terms proportional to $\tilde{\chi}_{nn^{\prime}}^{\lambda,\mathrm{PML},-}$, as given in Eqs. (71) and (72), respectively, which also involve coupling between the symmetrized modes. The gain components of the resonator are described by negative-energy contributions in the effective Hamiltonian. All the different bosonic modes couple to the QE with the coupling constants $g_{n,\lambda,X}$, as given in Eq. (75). Additional non-radiative decay or dephasing processes of the QE, here labeled by $\Gamma_{\mathrm{nr}}$, can be treated in the usual way by adding corresponding Lindblad terms to the master equation. (b) Scenario in which the electric/magnetic and gain/loss modal operators can be combined into a single operator per mode $\hat{a}_{n}$ (e.g., two spheres composed of passive, magnetodielectric media). In this case, all the gain, both from the PMLs and the resonator, is described by incoherent pumping terms proportional to $\tilde{\chi}_{nn^{\prime}}^{\mathrm{PML+G},-}$.
V Numerical Example: 1D Half-open Cavity
@@
Having presented the general formalism in the preceding sections, we now demonstrate the predictive capabilities of our approach in a simple, 1D example. As schematized in Fig. 3(a), we consider a half-open cavity comprising a homogeneous, diamond sample of relative permittivity $\varepsilon=5.7$ and length $L=200\,\textrm{nm}$, with a perfect electrical conductor boundary placed at $x=0$ (corresponding to a Dirichlet boundary condition imposed at this point). We assume that the diamond hosts a single QE at some location $x_{\mathrm{QE}}$ along the $x$ axis, featuring a pair of levels that are coupled by the electric dipole interaction with the cavity field. Such a QE could be realized in the form of a color center in the diamond lattice, and a variety of such optically active defects have been explored for quantum applications Aharonovich and Neu (2014). For the purpose of our numerical tests, we shall here consider a QE having the transition energy of a nitrogen-vacancy center, namely $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$ ($637\,\mathrm{nm}$), and a vertically oriented transition dipole moment ${\bm{d}}$ with controllable magnitude $|{\bm{d}}|$.
@@
A 1D formulation of mQED can be derived for systems that are invariant in two spatial dimensions (say, the $y$ and $z$ directions), and can be used to study the interaction of QEs with linearly polarized radiation propagating in the $x$ direction Gruner and Welsch (1996). In this case, we make the replacements $\hat{\bm{E}}(\bm{r})\rightarrow\hat{E}(x)=\hat{E}_{y}(x)$, ${\bm{d}}\rightarrow d=d_{y}$ and $\hat{\bm{f}}(\bm{r},\omega)\rightarrow\hat{f}(x,\omega)=\hat{f}_{y}(x,\omega)$. The $y$ component of the noise current density is then given by
@@
The formulae for the $P$-matrices all have the same form as in the 3D case, but with the replacements mentioned above and ×$\nabla\times\rightarrow\partial_{x}$:
$P_{nn^{\prime}}=\frac{\hbar\mu_{0}}{\pi A}\int_{0}^{\mathrm{\infty}}\mathrm{d}\omega\,\omega^{2}A_{n}(\omega)A_{n^{\prime}}^{*}(\omega)$ (85)
@@
where we have combined the electric and magnetic modal operators. The corresponding master equation then has the same form as in Eq. (81) with $\tilde{\chi}_{nn^{\prime}}^{\mathrm{PML+G},-}=0$ and $\tilde{P}_{nn^{\prime}}^{\mathrm{L}}=P_{nn^{\prime}}$. In the weak coupling or bad cavity limit, the electromagnetic modes can be traced out to obtain a reduced master equation for the QE only, as was done in Refs. Franke et al. (2019); Cirac (1992), with the QE decay rate
@@
For our calculations using Eq. (86), we retain the three lowest-order QNMs of the resonator which lie in spectral proximity to the QE transition $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$, with complex frequencies $\tilde{\omega}_{1}=(0.649-0.184\mathrm{i})\,\mathrm{eV}$, $\tilde{\omega}_{2}=(1.95-0.184\mathrm{i})\,\mathrm{eV}$ (i.e., resonant with the QE transition) and $\tilde{\omega}_{3}=(3.25-0.184\mathrm{i})\,\mathrm{eV}$, as shown in the upper panel of Fig. 3(b). These modes have very low quality factors of $Q_{1}=1.76$, $Q_{2}=5.28$ and $Q_{3}=8.80$. Furthermore, we assume that the QE is located at $x_{\mathrm{QE}}=66\,\mathrm{nm}$, where an electric-field antinode of the second-order QNM occurs [see Fig. 3(a)]. Under these conditions, whilst all three QNMs of the resonator offer polarization matching, only the second-order QNM allows optimal spectral and spatial matching with the QE transition. Fig. 3(b) presents our calculated Purcell factor, which bears the expected structure comprising three Lorentzian-like peaks at the spectral locations of the QNMs. In the weak coupling regime considered here, the decay rate can be derived semi-classically using Fermi’s golden rule and the Green’s function of the electric field Novotny and Hecht (2006), yielding (in 1D) $\Gamma^{\textrm{SC}}_{\mathrm{QE}}=\omega_{\mathrm{QE}}^{2}|d|^{2}\mathrm{Im}(G(x_{\mathrm{QE}},x_{\mathrm{QE}},\omega_{\mathrm{QE}}))/\hbar\varepsilon_{0}c^{2}A$. The predictions obtained using the fully quantum and semi-classical models are compared in the lower panel of Fig. 3(b), showing excellent agreement. Note that we have considered only a finite frequency interval $[\omega_{\mathrm{min}},\omega_{\mathrm{max}}]$ for the numerical integration in Eq. (85), with $\omega_{\mathrm{min}}=0.25\,\mathrm{eV}$ and $\omega_{\mathrm{max}}=4.0\,\mathrm{eV}$.
In the next step, we compare the electric field and QE dynamics in the strong coupling regime. In Ref. Dung, Knöll, and Welsch (2000); Scheel and Buhmann (2008), the spontaneous decay of QEs was studied within the framework of mQED. If the QE is prepared in the excited state |$\ket{\mathrm{e}}$ at time $t=0$, and the electromagnetic field in the vacuum state |$\ket{0}$, then the time-dependent state vector of the system can be written as |$\ket{\psi(t)}=C_{\mathrm{e}}(t)\mathrm{e}^{-\mathrm{i}\omega_{\mathrm{QE}}t}\ket{\mathrm{e}}\ket{0}+\int\mathrm{d}x\int\mathrm{d}\omega\,C_{\mathrm{g},\lambda}(x,\omega,t)\mathrm{e}^{-\mathrm{i}\omega t}\ket{\mathrm{g}}\hat{f}_{\lambda}^{\dagger}(x,\omega)\ket{0}$. From the Heisenberg equations of mQED, an integral equation for the time-dependent coefficient $C_{\mathrm{e}}(t)$ can be derived:
@@
This equation is derived from the exact mQED formulation and involves no approximations beyond the rotating-wave one for the interaction with the QE. We have solved the integral equation numerically, considering once more a QE tuned to the second-order QNM (i.e., $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$) and setting the dimensionless dipole moment to be $|d|/\sqrt{\hbar\varepsilon_{0}cA}=0.3$, which although unrealistically high, serves to demonstrate the validity of our theory in the strong coupling case. The result for the occupation probability of the excited state, $|C_{\mathrm{e}}(t)|^{2}$, can then be compared to the expectation value $\langle\sigma^{+}(t)\sigma^{-}(t)\rangle$, obtained from the quantum QNM master equation that we solve using the QuTiP library Lambert et al. (2024). The numerical data are plotted in Fig. 4(a), evidencing excellent agreement between the exact theory and the quantum QNM approach in capturing the temporal Rabi oscillations characteristic of this coupling regime. We have further plotted the expectation values of the QNM number operators $\langle\hat{a}^{\dagger}_{n}(t)\hat{a}_{n}(t)\rangle$; as expected, the symmetrized, second-order mode is dominant and undergoes coherent energy exchange with the QE, as reflected in the Rabi oscillations. Moreover, the amplitudes decrease in an approximately linear fashion on the logarithmic plot, with a gradient corresponding to the (equal) imaginary parts of the QNM frequencies $\Im(\tilde{\omega}_{n})=-0.184\,\mathrm{eV}$. After the first Rabi oscillation, the other two symmetrized modes show qualitatively similar oscillations as the second-order one, which might be explained by the contribution of the second-order QNM in the other symmetrized modes. Prior to this (i.e., during the first $\sim 10\,\mathrm{fs}$), their dynamics is more complicated and shows additional structure caused by the excitation of the first- and third-order QNMs.
From the exact mQED formulation, one can further derive an equation for the expectation value of the electric field intensity emitted via the spontaneous decay of the QE Dung, Knöll, and Welsch (2000):
@@
In Fig. 4(b), we compare the exact result for the normalized output intensity at $x=L$ with that obtained from the QNM master equation; once again, very good agreement can be seen. We observe two peaks of the intensity attributable to the Rabi oscillations of the second-order mode around $\sim 5\,\mathrm{fs}$ and $\sim 17\,\mathrm{fs}$. However, the first peak has additional substructure. Its maximum is split into a pair of maxima and the second derivative is changing sign along the drop of the peak. For comparison, we plot the intensity calculated from the quantum QNM model upon neglecting the first- and third-order modes (i.e., $|\tilde{E}^{\mathrm{sym}}_{2}(x)|^{2}\langle\hat{a}^{\dagger}_{2}(t)\hat{a}_{2}(t)\rangle$). From this, we can appreciate that the additional substructure originates from a multi-mode effect, as the intensity in this case only displays the two peaks from the Rabi oscillations.
@@
Figure 3: Numerical exploration of a 1D, half-open, diamond cavity in air, with an electric-dipole QE embedded therein. (a) Schematic of the cavity-QE system with characteristic parameters indicated. The dashed line shows the electric field intensity profile of the second-order QNM, to which the QE is resonantly coupled. (b) Complex frequencies $\tilde{\omega}_{n}$ of the three, lowest-order QNMs (upper panel) and the frequency-dependent Purcell factor of a QE with $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$ ($637\,\mathrm{nm}$), calculated in accordance with the quantum QNM and semi-classical theories (lower panel).

### v3

@@ IV Effective Lindblad Master Equation
Figure 2: Schematic representation of the quantum description of an electric/magnetic dipolar QE, modeled as a TLS, interacting with an electromagnetic resonator using PMLs. (a) Most general scenario, in which the resonator could have electric and magnetic gain components in distinct spatial or spectral regions (e.g., a pair of coupled, spherical resonators composed of different materials). In this case, every eigenmode of the resonator (both QNMs and PML modes) is described by four bosonic operators $\hat{a}_{n,\lambda,X}$ with $\lambda=\mathrm{e,m}$ and $X=\mathrm{L,G}$, corresponding to differently symmetrized modes according to Eqs. (45) and (46). The presence of gain/loss in the system necessitates a symmetrization and the consequent coupling between the symmetrized modes described by $\chi_{nn^{\prime}}^{\lambda,X,+}$ in the effective Hamiltonian of Eq. (62). The losses of the system are described by Lindblad dissipators proportional to $\chi_{nn^{\prime}}^{\lambda,\mathrm{L},-}$, as given in Eqs. (66), which also involve coupling between the symmetrized modes. The gain components of the resonator are described by negative-energy contributions in the effective Hamiltonian. All the different bosonic modes couple to the QE with the coupling constants $g_{n,\lambda,X}$, as given in Eq. (69). Additional non-radiative decay or dephasing processes of the QE, here labeled by $\Gamma_{\mathrm{nr}}$, can be treated in the usual way by adding corresponding Lindblad terms to the master equation. (b) Scenario in which the electric/magnetic and gain/loss modal operators can be combined intos a single operator per mode $\hat{a}_{n}$ (e.g., two spheres composed of passive, magnetodielectric media). In this case, the gain of the resonator is described by incoherent pumping terms proportional to $\tilde{\chi}_{nn^{\prime}}^{\mathrm{G},-}$.
V Numerical Examples
Having presented the general formalism in the preceding section, we now demonstrate the predictive capabilities of our approach by investigating the spontaneous emission of a QE in two structured, dielectric environments. In § V.1, we consider a 1D, half-open cavity due to its simple mode structure and analytical tractability. As the PMLs in 1D do not exhibit any gain components, we also briefly study a 3D spherical cavity in § V.2, where the proper elimination of the corresponding PML gain contributions enables accurate predictions to be ascertained.
V.1 1D Half-open Cavity
As schematized in Fig. 3(a), we consider a half-open cavity comprising a homogeneous, diamond sample of relative permittivity $\varepsilon=5.7$ and length $L=200\,\textrm{nm}$, with a perfect electrical conductor boundary placed at $x=0$ (corresponding to a Dirichlet boundary condition imposed at this point). In spite of its seeming simplicity, our choice of resonator provides a fairly stringent test of the theory, exhibiting QNMs with very low quality factors and spectrally overlapping resonance features in their coupling with QEs, a setting where phenomenological models based on normal modes can become inaccurate. Of course, our approach is expected to perform just as well with the more widely used, high-finesse Fabry-Pérot cavity designs of traditional quantum optics.
@@ V Numerical Examples
We assume that the diamond hosts a single QE at some location $x_{\mathrm{QE}}$ along the $x$ axis, featuring a pair of levels that are coupled by the electric dipole interaction with the cavity field. Such a QE could be realized in the form of a color center in the diamond lattice, and a variety of such optically active defects have been explored for quantum applications aharonovich2014diamond. For the purpose of our numerical tests, we shall here consider a QE having the transition energy of a nitrogen-vacancy center, namely $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$ ($637\,\mathrm{nm}$), and a vertically oriented transition dipole moment $\bm{d}$} with controllable magnitude $|{\bm{d}}|$.
@@
A 1D formulation of mQED can be derived for systems that are invariant in two spatial dimensions (say, the $y$ and $z$ directions), and can be used to study the interaction of QEs with linearly polarized radiation propagating in the $x$ direction gruner1996mqed. In this case, we make the replacements $\hat{\bm{E}}(\bm{r})\rightarrow\hat{E}(x)=\hat{E}_{y}(x)$, ${\bm{d}}\rightarrow d=d_{y}$ and $\hat{\bm{f}}(\bm{r},\omega)\rightarrow\hat{f}(x,\omega)=\hat{f}_{y}(x,\omega)$. The $y$ component of the noise current density is then given by
@@
The formulae for the $P$-matrices all have the same form as in the 3D case, but with the replacements mentioned above and $\nabla\times\rightarrow\partial_{x}$:
$P^{\mathrm{L}}_{nn^{\prime}}=\frac{\hbar\mu_{0}}{\pi A}\int_{0}^{\mathrm{\infty}}\mathrm{d}\omega\,\omega^{2}A_{n}(\omega)A_{n^{\prime}}^{*}(\omega)$ (79)
@@
where we have combined the electric and magnetic modal operators. The corresponding master equation then has the same form as in Eq. (65) with $\chi_{nn^{\prime}}^{\mathrm{G},-}=0$. In the weak coupling or bad cavity limit, the electromagnetic modes can be traced out to obtain a reduced master equation for the QE only, as was done in Refs. franke2019quantization; cirac1992interaction, with the QE decay rate
@@
For our calculations using Eq. (80), we retain the three lowest-order QNMs of the resonator which lie in spectral proximity to the QE transition $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$, with complex frequencies $\tilde{\omega}_{1}=(0.65-0.18\mathrm{i})\,\mathrm{eV}$, $\tilde{\omega}_{2}=(1.95-0.18\mathrm{i})\,\mathrm{eV}$ and $\tilde{\omega}_{3}=(3.25-0.18\mathrm{i})\,\mathrm{eV}$, as shown in the upper panel of Fig. 3(b). These modes have very low quality factors of $Q_{1}=1.76$, $Q_{2}=5.28$ and $Q_{3}=8.80$. Furthermore, we assume that the QE is located at $x_{\mathrm{QE}}=66\,\mathrm{nm}$, where an electric-field antinode of the second-order QNM occurs [see Fig. 3(a)]. Under these conditions, whilst all three QNMs of the resonator offer polarization matching, only the second-order QNM allows optimal spectral and spatial matching with the QE transition. Fig. 3(b) presents our calculated Purcell factor, which bears the expected structure comprising three Lorentzian-like peaks at the spectral locations of the QNMs. In the weak coupling regime considered here, the decay rate can be derived semi-classically using Fermi’s golden rule and the Green’s function of the electric field novotnyhechtbook, yielding (in 1D) $\Gamma^{\textrm{SC}}_{\mathrm{QE}}=\omega_{\mathrm{QE}}^{2}|d|^{2}\Im(G(x_{\mathrm{QE}},x_{\mathrm{QE}},\omega_{\mathrm{QE}}))/\hbar\varepsilon_{0}c^{2}A$. The predictions obtained using the fully quantum and semi-classical models are compared in the lower panel of Fig. 3(b), showing excellent agreement. Note that we have considered only a finite frequency interval $\omega_{\mathrm{min}},\omega_{\mathrm{max}}]$ for the numerical integration in Eq. (79), with $\omega_{\mathrm{min}}=0.25\,\mathrm{eV}$ and $\omega_{\mathrm{max}}=4.0\,\mathrm{eV}$.
[…]
@@
This equation is derived from the exact mQED formulation and involves no approximations beyond the rotating-wave one for the interaction with the QE. We have solved the integral equation numerically, considering a QE tuned to the second-order QNM (i.e., $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$) and setting the dimensionless dipole moment to be $|d|/\sqrt{\hbar\varepsilon_{0}cA}=0.3$, which although unrealistically high, serves to demonstrate the validity of our theory in the strong coupling case. The result for the occupation probability of the excited state, $|C_{\mathrm{e}}(t)|^{2}$, can then be compared to the expectation value $\langle\sigma^{+}(t)\sigma^{-}(t)\rangle$, obtained from the quantum QNM master equation that we solve using the QuTiP library lambert2024qutip. The numerical data are plotted in Fig. 4(a), evidencing excellent agreement between the exact theory and the quantum QNM approach in capturing the temporal Rabi oscillations characteristic of this coupling regime. We have further plotted the expectation values of the QNM number operators $\langle\hat{a}^{\dagger}_{n}(t)\hat{a}_{n}(t)\rangle$; as expected, the symmetrized, second-order mode is dominant and undergoes coherent energy exchange with the QE, as reflected in the Rabi oscillations. Moreover, the amplitudes decrease in an approximately linear fashion on the logarithmic plot, with a gradient corresponding to the (equal) imaginary parts of the QNM frequencies $\Im(\tilde{\omega}_{n})=-0.184\,\mathrm{eV}$. After the first Rabi oscillation, the other two symmetrized modes show qualitatively similar oscillations as the second-order one, which might be explained by the contribution of the second-order QNM in the other symmetrized modes. Prior to this (i.e., during the first $\sim 10\,\mathrm{fs}$), their dynamics is more complicated and shows additional structure caused by the excitation of the first- and third-order QNMs.
From the exact mQED formulation, one can further derive an equation for the expectation value of the electric field intensity emitted via the spontaneous decay of the QE dung2000mqed:
@@
In Fig. 4(b), we compare the exact result for the normalized output intensity at $x=L$ with that obtained from the QNM master equation; once again, very good agreement can be seen. We observe two peaks of the intensity attributable to the Rabi oscillations of the second-order mode around $\sim 5\,\mathrm{fs}$ and $\sim 17\,\mathrm{fs}$. However, the first peak has additional substructure. Its maximum is split into a pair of maxima and the second derivative is changing sign along the drop of the peak. For comparison, we plot the intensity calculated from the quantum QNM model upon neglecting the first- and third-order modes (i.e., $|\tilde{E}^{\mathrm{sym}}_{2}(x)|^{2}\langle\hat{a}^{\dagger}_{2}(t)\hat{a}_{2}(t)\rangle$). From this, we can appreciate that the additional substructure originates from a multimode effect, as the intensity in this case only displays the two peaks from the Rabi oscillations.
@@
Figure 3: Numerical exploration of a 1D, half-open, diamond cavity in air, with an electric-dipole QE embedded therein. (a) Schematic of the cavity-QE system with characteristic parameters indicated. The dashed line shows the electric field intensity profile of the second-order QNM. (b) Complex frequencies $\tilde{\omega}_{n}$ of the three, lowest-order QNMs (upper panel) and the frequency-dependent Purcell factor of the QE, calculated in accordance with the quantum QNM and semi-classical theories (lower panel).
@@
V.2 3D Spherical Cavity
To demonstrate the validity of our theory in the 3D case, we calculate the Purcell factor of a dipolar QE embedded in a spherical, dielectric resonator. Owing to the high degree of symmetry, both the QNMs and the $P$-matrix elements can be determined semi-analytically, just as in the 1D case, but with the crucial difference that in 3D, the PMLs include a gain contribution that must be correctly eliminated from the quantum dynamics, as discussed in § III. Specifically, we consider a silicon (Si) sphere with $\varepsilon_{\mathrm{Si}}=12$ and radius $R=850\,\mathrm{nm}$ in vacuum, as shown in Fig. 5(a). In the presence of ideal spherical symmetry, the modes can be classified as transverse magnetic and transverse electric modes, $\mathrm{TM}_{lmn}$ and $\mathrm{TE}_{lmn}$ respectively, with the angular quantum numbers $(l,m)$ ($l\geq 1$, $|m|\leq l$) and the radial quantum number $n$ ($n=1,2,\dotsstrattonbook$. Furthermore, TE and TM modes with different $l$ or $m$ do not pair in the $P$-matrix due to the orthogonality of the corresponding vector spherical harmonics, allowing for a separate treatment of each set of modes. As an illustration and for the sake of simplicity, we therefore consider three $\mathrm{TM}_{l=1}$ modes with radial quantum numbers $n=4,5,6$ (each with a three-fold degeneracy) and complex frequencies $\tilde{\omega}_{1m4}=(0.94-0.02\mathrm{i})\,\mathrm{eV}$, $\tilde{\omega}_{1m5}=(1.15-0.02\mathrm{i})\,\mathrm{eV}$ and $\tilde{\omega}_{1m6}=(1.37-0.02\mathrm{i})\,\mathrm{eV}$, which are plotted in the upper panel of Fig. 5(b). The contributions of these modes to the Purcell factor can be calculated as in Eq. (80), and as the analytic form of the Green’s function for such a system is decomposed into $\mathrm{TM}_{lm}$ and $\mathrm{TE}_{lm}$ parts li1994electromagnetic, we can compare the contribution of these modes directly with the exact $\mathrm{TM}_{l=1}$ contribution. The results for two different QE positions and orthogonal dipole orientations are shown in the lower panel of Fig. 5(b). We observe that the same resonances can have very different shapes for the two different dipoles, ranging from a Lorentzian-like peak to a strong Fano-like suppression of the decay rate, which the quantum QNM theory correctly reproduces to a high degree. The small deviation for the lowest-energy resonance likely stems from the contribution of other $\mathrm{TM}_{l=1}$ modes with lower radial quantum numbers $n$.
Figure 5: Numerical exploration of a 3D, spherical, Si cavity in air, with an electric-dipole QE embedded therein. (a) Schematic of the cavity-QE system with characteristic parameters indicated and two distinct coupling configurations, featuring a QE located at a radial distance of $r_{\bm{d}_{\parallel}}=380\,\mathrm{nm}$ or $r_{\bm{d}_{\perp}}=545\,\mathrm{nm}$ and with orthogonal dipole orientations. (b) Complex frequencies $\tilde{\omega}_{lmn}$ of the three considered $\mathrm{TM}_{l=1}$ QNMs (upper panel) and the frequency-dependent contributions of these modes to the Purcell factor of the QE in each configuration, calculated in accordance with the quantum QNM and semi-classical theories (lower panel).